OpenStudy (anonymous):

determine the slope and y intercept of the line represented by the given equation 2x-3y=15 what is the slope? what is the y intercept?

OpenStudy (anonymous):

2x=15+3y 3y= 2x-15 y= (2/3)x - 15/3 y= (2/3)x- 5 Slope: 2/3 y-intercept: (0,-5)

OpenStudy (anonymous):

i actually got this one right :)